OpenStudy (anonymous):

A parabola has a vertex at (-1, 0) and opens down. What is the equation of the parabola?

OpenStudy (anonymous):

@strawberryswing

OpenStudy (zale101):

y=a(x-h)^2+k is the vertex form of a parabola, where (h,k) represent the vertex (-1,0)

OpenStudy (zale101):

h=-1 k=0 and since there's no vertical stretching involved in this graph, then a will remain 1

OpenStudy (anonymous):

y ≤ -1 y ≥ -1 y ≤ 1 y ≥ 1

OpenStudy (anonymous):

@Zale101

OpenStudy (zale101):

your question says "the parabola opens down" y=a(x-h)^2+k a=-1 because it's a sad parabola 

OpenStudy (anonymous):

oh ok thanks :)